# Confronting Dark Matter Capture Rate with Continuous Gravitational Wave Probe of Local Neutron Stars

### v2

@@ I Introduction
It is nearly half a century ago, that the existence of dark matter (DM) has been confirmed from the observations of spiral galaxies. Later experiments like Planck Aghanim et al. (2020) also verified the presence of DM in the Universe with the observation of cosmic microwave background radiation. However, the nature of DM remains an enigma to be resolved as the particle physics of the Standard Model (SM) fails to provide a viable DM candidate. Despite its prevalence, the fundamental nature of DM remains unknown, fueling intense research efforts to detect and understand its properties. Among various DM searches, indirect detection of DM Slatyer (2018) has gained a lot of attention as the terrestrial direct detection experiments failed to detect any promising DM-nucleon scattering event. Indirect detection methods involve searching for the products resulting from annihilation or decay of DM particles ($\chi$), such as gamma rays, neutrinos, or cosmic rays, which can provide indirect evidence of its presence Gaskins (2016). This detection method has the potential to contribute to the ongoing quest to unravel the enigma of DM and shed light on the fundamental nature of the universe’s underlying structure.
@@
In this study, we undertake a comprehensive multi-wavelength approach, expanding our inquiry from gamma rays to neutrinos in the pursuit of detecting DM particles confined within NSs. This study aims to scrutinize the DM capture rate, particularly within nearby NSs located approximately 0.6 kpc from Earth, by probing the DM annihilation via long-lived mediator ($\phi$) with space-based or ground-based telescopes. The local NS population inferred from CGW emissions allows us to particularly investigate the cumulative population from around $\sim~{}10^{5}$ local NSs within 0.6 kpc Reed et al. (2021). Neutrino and gamma-ray searches offer unique advantages in the quest for DM. Both gamma rays and Neutrinos, being electrically neutral, travel directly from the source location unhindered by the Galactic or extra-galactic magnetic field. Neutrino signals, due to their weak interaction, travel unhindered from their sources without deflection or attenuation and can gain insights into dense sources, even those at cosmological distances, which other SM particles cannot reach. Telescopes like Fermi-LAT fer (2021); Ajello et al. (2021), LHAASO LHA (2021); Conceição (2023), CTA cta ; Donini et al. (2023), SWGO Conceição (2023), etc. search for sources of gamma rays from various celestial objects whereas IceCube Aartsen et al. (2020a), IceCube-Gen2 Abbasi et al. (2021), TRIDENT Ye et al. (2022), KM3NET Adrian-Martinez et al. (2016) experiments search for neutrino events from similar astrophysical bodies.
The paper is organized as follows. In Sec. II, we canvass the details of a few local NSs of our interest and nearby NS populations obtained from the CGW searches as primary sources of observation. In Sec. III, we brief on the theoretical framework of DM capture and annihilation within NS with long-lived mediators followed by a discussion on present and future experiments of significance to our analysis in Sec. IV. The bounds on DM scattering cross-sections obtained from the analysis with detector sensitivity for local NSs along with the NS population estimate of CGW searches are presented in Sec. V. We summarise the work with concluding remarks in Sec. VI.
@@ II.1 Observed nearby NSs
For this work, we consider the local NSs that lie within 0.6 kpc Vahdat et al. (2022) from Earth. We chose some nearby NSs from ref. Vahdat et al. (2022). We follow a few conditions to select our candidates:
  • Distance: There are nearby i.e. within 0.6 kpc from Earth.
  • Age: Old ($>$ Myr) so that DM annihilation rate and capture rate can reach equilibrium.
Table 1: Source Details: Column I: Name of our selected NS; Column II & III: R.A. and DEC; Column IV: Distance to the local NSs; Column V: Mass in $M_{\odot}$; Column VI: Radius ($R_{\rm NS}$) in km; Column VII: Age. We take the parameters of Column I-IV from ref. Vahdat et al. (2022). For mass and radius, we consider the general expected values for NS. For age estimation, we rely on the ATNF catalog atn .
@@ II.2 Local NS population predicted from CGW
In this section, we delve into the anticipated count of nearby NSs by exploring CGWs. The emission of gravitational waves from a rotating neutron star is contingent upon asymmetric deformations within its structure, manifesting as the star’s ellipticity ($\epsilon$) Zimmermann and Szedenits (1979); Lasky (2015). With an estimated age for the Milky Way of approximately $10^{10}$ years and a Galactic supernovae rate of 1 per century Diehl et al. (2006), it’s plausible that around $N_{0}\sim 10^{8}$ NSs have been generated within our Galaxy to date. However, only a fraction of these, primarily a few thousand, have been identified through electromagnetic searches, largely comprising radio pulsars Manchester et al. (2005). The detection of gravitational waves may facilitate the discovery of some of the remaining unidentified NSs and enable the investigation of their morphological distribution Caride et al. (2019); Abbott et al. (2019). A recent investigation by ref. Reed et al. (2021), conducted an extensive search for CGWs originating from unidentified Galactic NSs to establish constraints on their shapes. This study employed a straightforward model of the spatial and spin distributions of Galactic NSs to estimate the total count of NSs ($N_{\rm ns}$) within proximity to Earth, given a specific upper limit on $\epsilon$.
As detailed in Sec. II.1, our focus is solely on the local NSs within 0.6 kpc from Earth. Following the population model proposed by ref. Reed et al. (2021), the estimated number of NSs within a distance 0.6 kpc from Earth with $\epsilon~{}\sim~{}10^{-5}$ is found to be $N_{\rm ns}\simeq 10^{5}$. In subsequent sections, we elaborate on how such a large expected $N_{\rm ns}$ population within 0.6 kpc of Earth can impact our investigation.
@@ III Formulation
A) DM capture into NSs, contingent upon the local density and velocity distribution of DM, alongside the scattering cross-section of DM particles with neutrons.
B) The annihilation rate of DM particles ensnared gravitationally within NSs.
C) The emission of gamma-rays and neutrinos ensuing from the decay of long-lived mediators generated through DM annihilation.
@@ III.1 DM capture rate inside NSs
The “maximum capture rate” ($C_{\rm{max}}$) delineates the scenario where DM particles traversing through NSs become captured. An initial evaluation for $C_{\rm{max}}$ in a DM environment, characterized by parameters such as radius $R_{\rm NS}$, DM number density $n_{\chi}=\rho_{\chi}/m_{\chi}$, velocity dispersion ($\bar{v}$), and relative velocity ($v_{0}$), is given by Eq. 1. The escape velocity ($v_{\rm{esc}}$) is dictated by the mass $M_{\rm NS}$ and radius $R_{\rm NS}$ of the NS Leane and Smirnov (2021).
@@
Here, $v_{0}=\sqrt{8/(3\pi)}\bar{v}$, and $\bar{v}(r)=3/2v_{\rm{c}}(r)$, where $v_{\rm{c}}(r)=\sqrt{GM(r)/r}$ represents the circular velocity at any galactocentric distance $r$, with $G$ denoting the gravitational constant and $M(r)$ indicating the total Galactic mass encompassed within $r$. Regarding the DM density distribution, we adopt the Navarro-Frenk-White (NFW) density profile Navarro et al. (1996) with $\rho_{\chi(r=R_{\odot})}\equiv\rho_{0}=0.39$ GeV cm-3 Calore et al. (2022).
In a realistic setting, the perturbative approach must be accounted for, as not all DM particles passing through are captured. As mentioned, DM particles can undergo multiple scatterings, and when their velocity falls below the escape velocity of the NS, they become captured. The probability of experiencing $N$ scatterings before being trapped inside the NS is defined by Eq. 2, where $\tau=\frac{3}{2}\frac{\sigma_{\chi n}}{\sigma_{\rm{sat}}}$, with $\sigma_{\chi n}$ denoting the DM-nucleon scattering cross-section, $\sigma_{\rm{sat}}=\pi R_{\rm NS}^{2}/N_{n}$ representing the saturation cross-section, and $N_{n}=M_{\rm NS}/m_{n}$ indicating the number of nucleons in the target. The total capture rate ($C_{\rm{tot}}$) encompasses both single and multiple scatterings, as delineated in Eq. 3.
@@
The capture rate corresponding to $N$ scatterings, $C_{N}$, is delineated by Eq. 4 Bramante et al. (2017); Ilie et al. (2020). Here, $v_{N}$ is defined as $v_{N}=v_{\rm{esc}}(1-\beta_{+}/2)^{-N/2}$, where $\beta_{+}=4m_{\chi}m_{n}/(m_{\chi}+m_{n})^{2}$. For a large number of scatterings ($N\gg 1$), ($v_{N}^{2}-v_{\rm{esc}}^{2}$) predominates over $\bar{v}^{2}$, and $C_{N}$ approximates $\approx p_{N}C_{\rm{max}}$.
@@
III.2 DM annihilation to gamma rays and neutrinos
Ignoring the impact of evaporation (which is $\sim$ MeV Garani and Palomares-Ruiz (2022)), the total number of DM particles, represented as $N(t)$, accumulated within the core of the NS at time $t$, is governed by the DM number abundance in Eq. 5:
@@ III.2 DM annihilation to gamma rays and neutrinos
In this context, $C=\min[C_{\rm{tot}},C_{\rm{max}}]$ denotes the total capture rate under the condition that the perturbative estimation is deemed valid, while $C_{\rm ann}$ signifies the annihilation rate defined as $C_{\rm ann}=\frac{\langle\sigma v\rangle}{V_{\rm NS}}$; where $\langle\sigma v\rangle$ represents the annihilation cross-section for thermal DM and $V_{\rm NS}$ denotes the volume within the NS where annihilation occurs. If we disregard the influence of evaporation, the solution to Eq. 5 can be expressed as:
@@
The equilibrium time, $t_{\rm eq}~{}\equiv\frac{1}{\sqrt{C_{\rm ann}C}}$, represents the time required to reach equilibrium between DM capture rate and annihilation rate in the absence of evaporation. If equilibrium is achieved today (i.e. $t_{\rm NS}\geq t_{\rm eq}$), the total annihilation rate ($\Gamma_{\rm ann}$) is solely dependent on the DM capture rate, i.e. $\Gamma_{\rm ann}\rightarrow\frac{C}{2}$.
@@ III.3 Gamma-ray and neutrino spectrum: For DM annihilation via long-lived mediators
In this investigation, we delve into the emission of gamma rays and neutrinos resulting from the annihilation of DM. Typically, when DM particles undergo direct annihilation into quark-anti-quark pairs, they expect to generate detectable signals in the form of gamma rays and neutrinos as the final state products. However, when examining DM annihilation within celestial bodies such as NSs, there exists a possibility that if DM particles directly annihilate into SM final states, they could become confined within the NS, thereby inducing its heating. This heating phenomenon could potentially be observed using the James Webb Space Telescope (JWST), as recently discussed in reference Leane et al. (2021), although such observations pose significant challenges due to current instrumental limitations. Nevertheless, the chances of detecting a signal are enhanced when DM annihilation to SM states occurs via long-lived mediators ($\phi$) that can escape the NS’s surface, as proposed in various references Pospelov et al. (2008); Pospelov and Ritz (2009); Batell et al. (2010); Dedes et al. (2010); Fortes et al. (2016); Okawa et al. (2017); Yamamoto (2018); Holdom (1986a, b); Chen et al. (2009); Rothstein et al. (2009); Berlin et al. (2016); Cirelli et al. (2017, 2019). The decay products of these mediators may potentially yield observable signals in gamma-ray or neutrino telescopes.
Nonetheless, there remains a possibility that these mediators might interact with SM constituents within the NS before escaping, potentially resulting in a significant decrease in the observed flux of gamma rays and neutrinos. As illustrated in Leane et al. (2021); Bhattacharjee et al. (2023), it is possible to substantially mitigate this attenuation by carefully selecting appropriate model parameters. Hence, for the sake of simplicity and for ease of comparison with existing literature, we focus on the scenario where the long-lived mediator facilitates DM annihilation outside the NSs, such as $\chi\chi\rightarrow\phi\phi\rightarrow 4\gamma$ and $\chi\chi\rightarrow\phi\phi\rightarrow\nu\bar{\nu}\nu\bar{\nu}$. Additionally, we make the assumption that $m_{\phi}\ll m_{\chi}$, under which the provided formulas remain applicable even though this approximation removes the dependence on the additional parameter $m_{\phi}$.
The Eqs. 7 and 8 describe the differential flux of gamma rays and neutrinos reaching Earth as a result of DM annihilation via a long-lived mediator.
@@
$\frac{d\phi_{\nu}}{dE_{\nu}}=\frac{\Gamma_{\rm ann}}{4\,\pi\,d_{\rm NS}^{2}}\times\left(\frac{1}{3}\frac{dN_{\nu}}{dE_{\nu}}\right)\times P_{\rm surv},$ (8)
In this context, $d_{\rm NS}$ denotes the distance between Earth and the location of our selected NSs in sec. II.1. It also takes into account the contribution of the “survival probability” ($P_{\rm surv}$) for gamma rays and neutrinos reaching the detectors, which is nearly unity within the range of mediator decay lengths. Within the neutrino flux expression, there is a factor of $\frac{1}{3}$ to incorporate all neutrino flavors. In our scenarios, we are considering neutrinos arriving from distances of a few kpc, allowing us to confidently neglect the neutrino oscillation effects and proceed with the assumption that the proportion of neutrino flavors reaching Earth follows a ratio of $\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:1:1$.

### v3

# Confronting the dark matter capture rate with a continuous gravitational wave probe of local neutron stars

@@ I Introduction
Nearly half a century ago, the existence of dark matter (DM) was confirmed from the observations of spiral galaxies. Later experiments like Planck Aghanim et al. (2020) also verified the presence of DM in the Universe with the observation of cosmic microwave background radiation. However, the nature of DM remains an enigma that needs to be resolved as we are yet to detect a viable DM candidate. Despite its prevalence, the fundamental nature of DM remains unknown, fueling intense research efforts to detect and understand its properties. Among various DM searches, indirect detection of DM Slatyer (2018) has gained a lot of attention as the terrestrial direct detection experiments failed to detect any promising DM-nucleon scattering event. Indirect detection methods involve searching for the products resulting from annihilation or decay of DM particles ($\chi$), such as gamma rays, neutrinos, or cosmic rays, which can provide indirect evidence of its presence Gaskins (2016). This detection method has the potential to contribute to the ongoing quest to unravel the enigma of DM and shed light on the fundamental nature of the Universe’s underlying structure.
@@
In this study, we undertake a comprehensive multiwavelength approach, expanding our inquiry from gamma rays to neutrinos in the pursuit of detecting DM particles confined within NSs. This study aims to scrutinize the DM capture rate, particularly within nearby NSs located approximately 0.6 kpc from Earth, by probing the DM annihilation via long-lived mediator ($\phi$) with space-based or ground-based telescopes. The local NS population inferred from CGW emissions allows us to particularly investigate the cumulative population from around $\sim~{}10^{5}$ local NSs within 0.6 kpc Reed et al. (2021). Neutrino and gamma-ray searches offer unique advantages in the quest for DM. Both gamma rays and Neutrinos, being electrically neutral, travel directly from the source location unhindered by the Galactic or extragalactic magnetic field. Neutrino signals, due to their weak interaction, travel unhindered from their sources without deflection or attenuation and can gain insights into dense sources, even those at cosmological distances, which other SM particles cannot reach. Telescopes like Fermi-LAT fer (2021); Ajello et al. (2021), LHAASO LHA (2021); Conceição (2023), CTA cta ; Donini et al. (2023), SWGO Conceição (2023), etc. search for sources of gamma rays from various celestial objects whereas IceCube Aartsen et al. (2020a), IceCube Gen 2 Abbasi et al. (2021), TRIDENT Ye et al. (2022), KM3NET Adrian-Martinez et al. (2016) experiments search for neutrino events from similar astrophysical bodies.
The paper is organized as follows. In Sec. II, we canvass the details of a few local NSs of our interest and nearby NS populations obtained from the CGW searches as primary sources of observation. In Sec. III, we brief on the theoretical framework of DM capture and annihilation within NS with long-lived mediators followed by a discussion on present and future experiments of significance to our analysis in Sec. IV. The bounds on DM scattering cross sections obtained from the analysis with detector sensitivity for local NSs along with the NS population estimate of CGW searches are presented in Sec. V. We summarise the work with concluding remarks in Sec. VI and a discussion in Sec. VII addressing possible improvements of the results when corrections are implemented on present formalism.
@@ II Source Selection
In this section, we briefly discuss the local NSs of our interest and introduce the local NS population with continuous gravitational wave searches.
@@ II.1 Observed nearby NSs
In this study, we focus on nearby local NSs. We use two criteria to select our candidates: (i) distance from Earth, and (ii) their age. Following Vahdat et al.(2021) Vahdat et al. (2022), which provides a list of local NSs, we constrain our selection to those within 0.6 kpc and older than $10^{6}$ years, resulting in a selection of 8 nearby NSs. For the age criterion, we use the data from the ATNF catalog atn . The positions and characteristics of the selected sources are presented in Table 1.
[…]
Table 1: Source Details: Column I: Name of our selected NS; Column II and III: R.A. and DEC in degree; Column IV: Distance ($d_{\rm NS}$) to the local NSs in kpc; Column V: Mass ($M_{\rm NS}$) in $M_{\odot}$; Column VI: Radius ($R_{\rm NS}$) in km; Column VII: Age in year. We take the parameters of column I-IV from Ref. Vahdat et al. (2022). For mass and radius, we consider the general expected values for NS. For age estimation, we rely on the ATNF catalog atn .
@@ II.2 Local NS population predicted from CGW
Indirect detection of DM and study of its properties from NSs are based on the gamma-ray and neutrino signal from the sources, focusing mainly on the distribution of the galactic NS population. However, only a very few local NSs are detected making it difficult to constrain the DM-nucleon scattering. As a result, local NSs are generally not considered very significant for the search of DM even though they are expected to be one of the most promising celestial objects to capture DM. This motivates us to look beyond the detection of local NSs with complementary experiments and continuous gravitational wave (CGW) searches provide us with a new window of opportunity. Distinguishing the CGW from local NSs from the background CGW not only leads to the detection of unknown NSs but also helps to predict the local NS distribution in the neighborhood of Earth. Hence, with the help of the local NS population from CGW searches, we expect to obtain a significant improvement in the existing limits on DM capture rate by NSs, especially from the observational perspective.
In this section, we delve into the anticipated count of nearby NSs by exploring CGWs. With an estimated age for the Milky Way of approximately $10^{10}$ years and a Galactic supernovae rate of 1 per century Diehl et al. (2006), it is plausible that around $N_{0}\sim 10^{8}$ NSs have been generated within our Galaxy to date. However, only a fraction of these, primarily a few thousand, have been identified through electromagnetic searches, largely comprising radio pulsars Manchester et al. (2005). Spinning NSs with deformity in their shapes are promising sources of gravitational waves. The detection of gravitational waves originating from NSs may facilitate the discovery of some of the remaining unidentified NSs and enable the investigation of their morphological distribution Caride et al. (2019); Abbott et al. (2019). A recent investigation by Ref. Reed et al. (2021), conducted an extensive search for CGWs originating from unidentified Galactic NSs to establish constraints on their shapes. This study employed a straightforward model of the spatial and spin distributions of Galactic NSs to estimate the total count of NSs ($N_{\rm NS}$) within proximity to Earth, given a specific upper limit on $\epsilon$.
The emission of gravitational waves from a rotating NS is contingent upon asymmetric deformations within its structure, manifesting as the star’s ellipticity ($\epsilon$) Zimmermann and Szedenits (1979); Lasky (2015), defined as
$\epsilon=\sqrt{\frac{8\pi}{15}}\frac{Q_{22}}{I_{zz}}\,,$ (1)
where quadrupole moment due to deformation is $Q_{22}$, and $I_{zz}$ is the principal moments of inertia along $z$ direction. Considering a rotating NS at a distance $d_{\rm NS}$ and spin frequency $\nu$. The corresponding strain amplitude $h_{0}$ of the NS is expressed as
$h_{0}=\frac{4\pi^{2}G_{\rm N}}{c^{4}}\frac{I_{zz}f_{\rm GW}^{2}}{d_{\rm NS}}\epsilon\ ,$ (2)
[…]
@@ III Formulation
(A) DM capture into NSs, contingent upon the local density and velocity distribution of DM, alongside the scattering cross section of DM particles with neutrons.
(B) The annihilation rate of DM particles ensnared gravitationally within NSs.
(C) The emission of gamma-rays and neutrinos ensuing from the decay of long-lived mediators generated through DM annihilation.
@@ III.1 DM capture rate inside NSs
The “maximum capture rate” ($C_{\rm{max}}$) delineates the scenario where DM particles traversing through NSs become captured. An initial evaluation for $C_{\rm{max}}$ in a DM environment, characterized by parameters such as radius $R_{\rm NS}$, DM number density $n_{\chi}=\rho_{\chi}/m_{\chi}$, velocity dispersion ($\bar{v}$), and relative velocity ($v_{0}$), is given by Eq. (3). The NS absorbs the kinetic energy of DM when DM is accelerated to relativistic speeds by its intense gravitational potential. To properly account for the gravitational focusing effect, the escape velocity ($v_{\rm{esc}}$) is modified and expressed as $\sqrt{2\chi}$, where $\chi=1-\sqrt{1-\frac{2G_{\rm N}M_{\rm NS}}{R_{\rm NS}}}$ Leane and Smirnov (2021); Bell et al. (2020); Leane and Tong (2024).
@@
Here, $v_{0}=\sqrt{8/(3\pi)}\bar{v}$, and $\bar{v}(r)=3/2v_{\rm{c}}(r)$, where $v_{\rm{c}}(r)=\sqrt{G_{\rm N}M(r)/r}$ represents the circular velocity at any galactocentric distance $r$, with $M(r)$ indicating the total Galactic mass encompassed within $r$. Regarding the DM density distribution, we adopt the Navarro-Frenk-White (NFW) density profile Navarro et al. (1996), as it is widely used in the literature and provides a well-motivated framework consistent with both observational data and the theoretical understanding of DM halos, with $\rho_{\chi(r=R_{\odot})}\equiv\rho_{0}=0.39$ GeV cm-3 Calore et al. (2022).
In a realistic setting, the perturbative approach must be taken into account, as not all DM particles passing through are captured. As mentioned, DM particles can undergo multiple scatterings, and when their velocity falls below the escape velocity of the NS, they become captured. The probability of experiencing $N$ scatterings before being trapped inside the NS is defined by Eq. (4), where $\tau=\frac{3}{2}\frac{\sigma_{\chi n}}{\sigma_{\rm{sat}}}$, with $\sigma_{\chi n}$ denoting the DM-nucleon scattering cross section, $\sigma_{\rm{sat}}=\pi R_{\rm NS}^{2}/N_{n}\simeq 6.58\times 10^{-45}$ cm2 Bramante et al. (2017), representing the saturation cross section, and $N_{n}=M_{\rm NS}/m_{n}$ indicating the number of nucleons in the target. As highlighted in recent works Bell et al. (2020, 2021d); Anzuini et al. (2021), $N_{n}$ and $v_{\rm{esc}}$ might not be constant and vary across the NS interior due to strong gravitational effects and varying densities. However, in this work, we have adopted a simplified model for direct comparisons and benchmarking of our results within a familiar context Leane and Tong (2024); Acevedo et al. (2024); Leane et al. (2021); Bose et al. (2022), we assume $N_{n}$ and $v_{\rm{esc}}$ as constants within the NS interior.
The total capture rate ($C_{\rm{tot}}$) encompasses both single and multiple scatterings, as delineated in Eq. (5).
@@
The capture rate corresponding to $N$ scatterings, $C_{N}$, is delineated by Eq. (6) Bramante et al. (2017); Ilie et al. (2020). Here, $v_{N}$ is defined as $v_{N}=v_{\rm{esc}}(1-\beta_{+}/2)^{-N/2}$, where $\beta_{+}=4m_{\chi}m_{n}/(m_{\chi}+m_{n})^{2}$. For a large number of scatterings ($N\gg 1$), ($v_{N}^{2}-v_{\rm{esc}}^{2}$) predominates over $\bar{v}^{2}$, and $C_{N}$ approximates $\approx p_{N}C_{\rm{max}}$.
@@
We emphasize that for $m_{\chi}<1$ GeV, the accretion rate no longer scales as $1/m_{\chi}$ but is instead suppressed due to Pauli blocking (PB). We scale the $\sigma_{\rm{sat}}$ value for $m_{\chi}<1$ GeV by following the Ref. Baryakhtar et al. (2017). As reported in some recent studies Bell et al. (2020); Garani et al. (2019), we analyze the changes in $C_{\rm{tot}}$ for $m_{\chi}<1$ GeV in comparison to $C_{\rm{tot}}$ at $m_{\chi}=1$ GeV. Following the Pauli blocking suppression reported in Ref. Bell et al. (2020) with respect to $C_{\rm{tot}}$ at $m_{\chi}=1$ GeV, we define a scaling factor $f_{\rm PB}$ which is $\sim 1.6\times m_{\chi}^{1.2}$ valid up to $m_{\chi}<1$ GeV  [^3] It is important to note that the $\sigma_{\rm{sat}}$ values for the selected sources are comparable to the geometric limits used in Ref. Bell et al. (2020). As a result, our scaling formula provides a reasonably accurate estimate of the expected Pauli blocking effect, avoiding significant overestimation or underestimation. By applying this scaling factor to Eq. (5), we determine the capture rate under Pauli blocking suppression, $C_{\rm{PB}}(r)$, for $m_{\chi}<1$ GeV, given by,
$C_{\rm{PB}}(r)=C_{\rm{tot}}(r)\times f_{\rm PB}$ (7)
III.2 DM annihilation rate
In this section, we discuss the DM annihilation rate resulting from accumulated DM inside NS. We want to point out that for this study we ignore the impact of DM evaporation, effective for $m_{DM}\leq$ 10 eV for a Gyr old NSBell et al. (2021e, 2024). Therefore, in the absence of evaporation, the total number of DM particles, represented as $N(t)$, accumulated within the core of the NS at time $t$, is governed by the evolution of DM number abundance in Eq. (8):
@@ III.2 DM annihilation rate
In this context, $C=\min[C_{\rm{tot}},C_{\rm{max}}]$ denotes the total capture rate under the condition that the perturbative estimation is deemed valid, while $C_{\rm ann}$ signifies the annihilation rate defined as $C_{\rm ann}=\frac{\langle\sigma v\rangle}{V_{\rm NS}}$; where $\langle\sigma v\rangle$ represents the annihilation cross section for thermal DM and $V_{\rm NS}$ denotes the volume within the NS where annihilation occurs. As we disregard the influence of evaporation, the solution to Eq. (8) can be expressed as:
@@
The equilibrium time, $t_{\rm eq}~{}\equiv\frac{1}{\sqrt{C_{\rm ann}C}}$, represents the time required to reach equilibrium between the DM capture rate and the annihilation rate in the absence of evaporation. If equilibrium is achieved today (i.e., $t_{\rm NS}\geq t_{\rm eq}$), the total annihilation rate ($\Gamma_{\rm ann}$) is solely dependent on the DM capture rate, i.e., $\Gamma_{\rm ann}\rightarrow\frac{C}{2}$. As mentioned in Ref. Bell et al. (2024), the equilibrium age of an NS is under one year for vector interactions and approximately $10^{4}$ years for scalar interactions. Given that all our sources are older than $10^{6}$ years, we can conclude that the equilibrium between capture and annihilation rates has been established.
@@ III.3 Gamma-ray and neutrino spectrum: For DM annihilation via long-lived mediators
In this investigation, we aim to probe the gamma-ray and the neutrino emission resulting from the DM annihilation inside NS. Typically, when DM particles undergo direct annihilation into quark-antiquark pairs, they expect to generate detectable signals in the form of gamma rays, neutrinos, or any other form of SM products as final states. However, when examining DM annihilation within dense celestial bodies such as NSs, there exists a possibility that if DM particles directly annihilate into SM final states, they might become confined within the NS, thereby inducing its heating. This heating phenomenon could potentially be observed using the James Webb Space Telescope (JWST), as recently discussed in reference Leane et al. (2021), although such observations pose significant challenges due to current instrumental limitations. Nevertheless, the chances of detecting a signal are enhanced when DM annihilation to SM states occurs via long-lived mediators ($\phi$) that can escape the NS’s surface, as proposed in various references Pospelov et al. (2008); Pospelov and Ritz (2009); Batell et al. (2010); Dedes et al. (2010); Fortes et al. (2016); Okawa et al. (2017); Yamamoto (2018); Holdom (1986a, b); Chen et al. (2009); Rothstein et al. (2009); Berlin et al. (2016); Cirelli et al. (2017, 2019). The decay products of these mediators may potentially yield observable signals in gamma-ray or neutrino telescopes.
Nonetheless, there remains a possibility that these mediators might interact with SM constituents within the NS before escaping, potentially resulting in a significant decrease in the observed flux of gamma rays and neutrinos. As illustrated in Leane et al. (2021); Bhattacharjee et al. (2023), it is possible to substantially mitigate this attenuation by carefully selecting appropriate model parameters. Hence, for the sake of simplicity and ease of comparison with existing literature, we focus on the scenario where the long-lived mediator facilitates DM annihilation outside the NSs, such as $\chi\chi\rightarrow\phi\phi\rightarrow 4\gamma$ and $\chi\chi\rightarrow\phi\phi\rightarrow\nu\bar{\nu}\nu\bar{\nu}$. Additionally, we make the assumption that $m_{\phi}\ll m_{\chi}$, under which the provided formulas remain applicable even though this approximation removes the dependence on the additional parameter $m_{\phi}$. It is to be noted that, we consider DM annihilates into photon or neutrino only with branching ratio unity, and has no other annihilation channel. This corresponds to the maximum flux obtained from DM into the specific annihilation channel. In general, DM can have many annihilation channels depending on the mass. With the specific choice of models where DM is neutrinophilic, and the mediator $\phi$ is coupled to neutrino only, this kind of DM annihilation scenario can be achieved. However, constructing a complete model for DM is beyond the scope of this work.
The Eqs. (10) and (11) describe the differential flux of gamma rays and neutrinos reaching Earth as a result of DM annihilation via a long-lived mediator.
@@
$\frac{d\phi_{\nu}}{dE_{\nu}}=\frac{\Gamma_{\rm ann}}{4\,\pi\,d_{\rm NS}^{2}}\times\left(\frac{dN_{\nu}}{dE_{\nu}}\right)\times P_{\rm surv},$ (11)
In this context, we take into account the contribution of the “survival probability” ($P_{\rm surv}$) for gamma rays and neutrinos reaching the detectors, which is nearly unity within the range of mediator decay lengths. In our scenarios, we are considering neutrinos arriving from distances of a few kpc, allowing us to confidently neglect the neutrino oscillation effects and proceed with the assumption that the proportion of neutrino flavors reaching Earth follows a ratio of $\nu_{e}:\nu_{\mu}:\nu_{\tau}=1:1:1$.